Circular No. 6133

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

POSSIBLE NOVA IN AQUILA S. Nakano, Sumoto, Japan, reports the discovery of a possible nova

by Kesao Takamizawa, Saku-machi, Nagano-ken, on two T-Max 400 films taken Feb. 7.839 and 7.842 UT using a 10-cm f/4.0 lens. Y. Kushida, Yatsugatake South Base Observatory, provides the following position: α = 19^{h}05^{m}26^{s}.57, δ = −1^{o}42′03′′.5. Takamizawa notes that there is no image brighter than mag 12–14 at this position on films taken on 1994 May 12, June 14, and Oct. 23; there is a nearby star of mag 12.1. Available magnitude estimates, the last two being visual estimates supplied by E. Waagen, AAVSO: Feb. 7.839, 8.1 (Takamizawa); 8.854, 8.4 (Takamizawa); 8.873, 7.5: (Kushida); 9.435, 8.2 (C. Scovil, Stamford, CT); 9.531, 8.5 (P. Collins, Scottsdale, AZ). Collins notes that the star has a reddish color.

SUPERNOVA 1995C IN ANONYMOUS GALAXY Robert H. McNaught, Anglo-Australian Observatory, reports his dis-

covery of an apparent supernova (mag ∼ 17.5) on a U.K. Schmidt Telescope (UKST) R plate taken by K. S. Russell on Feb. 8.74 UT. SN 1995C is located at α = 12^{h}04^{m}02^{s}.52, δ = −31^{o}36′02′′.3 (equinox 2000.0, uncertainty 0′′.2 in α and 0′′.4 in δ), which is 14′′.7 east and 8′′.5 south of the galaxy’s center. The object does not appear on an ESO R survey plate or on a UKST R plate taken on 1993 May 23. Confirmation was made via an unfiltered CCD image taken Feb. 9.56 by McNaught and D. J. Asher with the 1.0-m telescope at Siding Spring Observatory. A nearby star of mag R ∼ 18 has position end figures 03^{s}.07, 26′′.0.

COMET 19P/BORRELLY Total visual magnitude estimates: Jan. 27.89 UT, 9.7 (G. A. Milani,

Cima Ekar, Italy, 20×80 binoculars); 30.24, 10.1 (C. S. Morris, Whitaker Peak, CA, 0.26-m reflector); Feb. 3.15, 10.4 (A. Hale, near Chaparral, NM, 0.2-m reflector); 8.10, 10.7 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

COMET C/1994 T1 (MACHHOLZ) Total visual magnitude estimates by Morris at Whitaker Peak: Jan.

2.21 UT, 11.8; Feb. 1.18, 13.3.

1995 February 9 Daniel W. E. Green